# Cooperative Nonlinear Guidance Strategies for Guaranteed Pursuit-Evasion

### v2

@@ IV Simulations
(b) Sliding manifold and the error.
@@
(b) Sliding manifold and the error.
(c) Steering controls of the evader and the defender.
(d) Joint effort of the evader-defender team and the pursuer’s maneuver.
@@
(b) Sliding manifold and the error.
(c) Steering controls of the defender and the pursuer.
@@
(b) Sliding manifold and the error.
(c) Steering controls of the defender and the pursuer.
@@
(b) Sliding manifold and the error.
@@
We now demonstrate the performance of the proposed nonlinear guidance strategy using simulation for various three-body engagement scenarios. We assume $v_{\mathrm{P}}=v_{\mathrm{D}}=200$ m/s, whereas $v_{\mathrm{E}}=100$ m/s. In each case, we place the evader at the origin of the inertial place at the beginning of the engagement. The initial positions of the agents in the trajectory plots are denoted using square markers, whereas an intercept is represented using an asterisk. In the following plots, the pursuer’s trajectories and inputs are shown in red, while those of the evader and defender are shown in blue and green, respectively. In our simulations, the pursuer and the defender possess the same maneuverability capabilities, limited to ±$\pm 20$ g. On the other hand, the evader can only apply a maximum steering control of $10$ g in either direction. Here g$=9.81$ m/s2 is the acceleration due to gravity. Thus, the pursuer and the defender are evenly matched in terms of both speed and acceleration capabilities. The design parameters are chosen to be $t_{1}=3$ s, $t^{\star}=6$ s, $k_{1}=6$, $k_{2}=3$ and $\mathcal{K}=5$. Such choices of $t_{1}$ and $t^{\star}$ ensure that the sliding manifold converges to zero within 3 s, whereas the error vanishes within 6 s.
Consider a scenario where the pursuer is using a guidance strategy that is a function of the pursuer-evader LOS rate ($\dot{\lambda}_{\mathrm{EP}}$), for example, proportional-navigation guidance. For the pursuer, such a strategy may also be optimal in certain cases as it aims to arrive on the collision course with minimum effort. The guidance command for the pursuer is given by $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, where $\mathscr{N}=3$ is a constant. This situation is shown in Figure 3. The defender and pursuer are initially located $400$ m and $5000$ m radially apart from the evader, with respective LOS angles of $-45^{\circ}$ and $0^{\circ}$. The initial heading angles of the agents are $\gamma_{\mathrm{E}}=45^{\circ}$, $\gamma_{\mathrm{D}}=0^{\circ}$, and $\gamma_{\mathrm{P}}=180^{\circ}$. Under these settings, the initial value of the angle $\chi$ is $228^{\circ}$. It can be observed from Figure 3(a) that based on different values of the desired $\chi^{\star}$, the agents maneuver differently. However, the pursuer is always intercepted before it can capture the evader. Figure 3(b) depicts the sliding manifold and the error profiles of various desired angles $\chi^{\star}$, which evidences that $\underaccent{\bar}{\mathcal{S}}$ converges within 3 s and $\beta$ nullifies within 6 s, as expected. Therefore, regardless of the three-body initial engagement geometry, the evader-defender team maneuvers to ensure the pursuer’s capture. Figures 3(c) and 3(d) illustrate the various control efforts of the agents. It is important to note that the efforts of the evader and the defender (individually and jointly) have small magnitudes in the endgame.
Since the proposed defense strategy is independent of the pursuer’s maneuver, the defender is able to capture the pursuer even if it uses a different guidance law. Figure 4 portrays various cases when the pursuer uses proportional-navigation guidance (PNG), pure pursuit guidance (PPG), and deviated pursuit guidance (DPG) to capture the evader. The guidance command for PNG, PPG, and DPG are chosen to be $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}\right)$, and $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}-\delta\right)$, respectively with $\mathscr{N}=3$ and $\delta=20^{\circ}$. For a fixed angle, $\chi^{\star}=180^{\circ}$, the defender always intercepts the pursuer midway and safeguards the evader. In Figure 4, the initial conditions are kept the same as that in the previous case, except $\chi^{\star}=180^{\circ}$. We also observe similar behaviors of the sliding manifold, error, and the agents’ control efforts.
We now consider cases when the evader is stationary, so the pursuer can head directly toward it with an optimal effort once any heading angle errors vanish. In Figure 5, the pursuer starts at a distance of 5000 m from the evader with a LOS angle of $2^{\circ}$. The defender, on the other hand, starts at three different positions in three different cases. With respect to the evader, the defender is at a relative separation of $400$ m, $1500$ m, and $2000$ m with LOS angles of $-45^{\circ}$, $-10^{\circ}$, and $10^{\circ}$, respectively. The defender aims to attain $\chi^{\star}=180^{\circ}$ in each case to arrive directly between the evader and the pursuer such that the latter’s capture can be guaranteed. We notice that even if the evader does not maneuver, the defender maneuvers accordingly to satisfy the specific geometrical conditions for a constant $\chi^{\star}$, thereby guaranteeing pursuit-evasion. On the other hand, the situation in Figure 6 assumes a fixed position of the defender ($r_{\mathrm{DE}}=400$ m, $\lambda_{\mathrm{DE}}=-45^{\circ}$), whereas the pursuer is located at different positions in different cases. The initial configurations of the pursuer relative to the evader are $r_{\mathrm{EP}}=5000$ m with LOS angles of $-5^{\circ},0^{\circ},$ and $2^{\circ}$, respectively. Once again, we observe that the fixed angle $\chi^{\star}$ is achieved by the defender in $t^{\star}=6$ s and maintained thereafter to ensure the pursuer’s capture. We now demonstrate the performance of the proposed strategy when the maximum steering controls of the defender and pursuer are chosen to be $8$ g, whereas the maximum steering control input for the evader is chosen to be $4$ g. While the other engagement settings are kept the same as for $\chi^{\star}=180^{\circ}$, and we depict the performance under reduced maximum capability through Figure 7(d). It can be observed that the defender intercepts the pursuer by following a similar behavior as when maximum accelerations are higher. This further bolsters the arguments presented in this work that pursuit-evasion is guaranteed regardless of the three-body initial engagement geometry and the knowledge of the pursuer’s strategy.
V Conclusions
@@
In this paper, we have introduced a geometric approach that addresses the challenge of pursuit-evasion scenarios involving three agents with arbitrary initial geometries. Our proposed solution, the evader-defender cooperative guidance strategy, offers an effective means to guarantee pursuit-evasion under diverse conditions. Specifically, it guarantees that the defender reaches a specific angle precisely within a predefined time, irrespective of the initial engagement geometry, thus intercepting the pursuer before it can capture the evader. A distinguishing feature of our approach is its adaptability to scenarios characterized by nonlinear dynamics, non-holonomic constraints, and large heading angle errors, which are often encountered in practical motion control situations, including aircraft defense. This paper not only presents a robust and practical solution for guaranteed pursuit-evasion problems but also lays the foundation for identifying crucial conditions and configurations that facilitate successful evasion or capture. Analyzing geometrical solutions in three-dimensional settings could be an interesting future research.

### v3

@@ 4.1 Maneuver-level cooperation between evader and defender
(b) Sliding manifold and error.
@@
Since the proposed defense strategy is independent of the pursuer’s maneuver, the defender is able to capture the pursuer even if it uses a different guidance law. Figure 4 portrays various cases when the pursuer uses proportional-navigation guidance (PNG), pure pursuit guidance (PPG), and deviated pursuit guidance (DPG) to capture the evader. The guidance command for PNG, PPG, and DPG are chosen to be $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}\right)$, and $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}-\delta\right)$, respectively with $\mathscr{N}=3$ and $\delta=20^{\circ}$. For a fixed angle, $\chi^{\star}=180^{\circ}$, the defender always intercepts the pursuer midway and safeguards the evader. In Figure 4, the initial conditions are kept the same as those in the previous case, except $\chi^{\star}=180^{\circ}$. We also observe similar behaviors of the sliding manifold, error, and the agents’ control efforts.
@@
(b) Sliding manifold and error.
(c) Steering controls $a_{\rm D}$ and $a_{\rm E}$.
(d) $\mathcal{U}$ and $a_{\rm P}$.
@@
We now consider cases when the evader is stationary, so the pursuer can head directly toward it with an optimal effort once any heading angle errors vanish. In Figure 5, the pursuer starts at a distance of 5000 m from the evader with a LOS angle of $2^{\circ}$. The defender, on the other hand, starts at three different positions in three different cases. With respect to the evader, the defender is at a relative separation of $400$ m, $1500$ m, and $2000$ m with LOS angles of $45^{\circ}$, $10^{\circ}$, and $10^{\circ}$, respectively. The defender aims to attain $\chi^{\star}=180^{\circ}$ in each case to arrive directly between the evader and the pursuer such that the latter’s capture can be guaranteed. We notice that even if the evader does not maneuver, the defender maneuvers accordingly to satisfy the specific geometrical conditions for a constant $\chi^{\star}$, thereby guaranteeing pursuit-evasion. On the other hand, the situation in Figure 6 assumes a fixed position of the defender ($r_{\mathrm{DE}}=400$ m, $\lambda_{\mathrm{DE}}=-45^{\circ}$), whereas the pursuer is located at different positions in different cases. The initial configurations of the pursuer relative to the evader are $r_{\mathrm{EP}}=5000$ m with LOS angles of $5^{\circ},0^{\circ},$ and $2^{\circ}$, respectively. Once again, we observe that the fixed angle $\chi^{\star}$ is achieved by the defender in $t^{\star}=6$ s and maintained thereafter to ensure the pursuer’s capture.
@@
(b) Sliding manifold and error.
(c) Steering controls $a_{\rm D}$ and $a_{\rm P}$.
@@
(b) Sliding manifold and error.
(c) Steering controls of defender and pursuer.
@@
We now demonstrate the performance of the proposed strategy when the maximum steering controls of the defender and pursuer are chosen to be $8$ g, whereas the maximum steering control input for the evader is chosen to be $4$ g. While the other engagement settings are kept the same as for $\chi^{\star}=180^{\circ}$, we depict the performance under reduced maximum capability through Figure 7(d). It can be observed that the defender intercepts the pursuer by following a similar behavior as when maximum accelerations are higher. This further bolsters the arguments presented in this work that pursuit-evasion is guaranteed regardless of the three-body initial engagement geometry and the knowledge of the pursuer’s strategy.
@@
(b) Sliding manifold and error.
@@
(a) Trajectories.
(b) Sliding manifold and error.
(c) Steering controls $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$.
(d) $a_{\mathrm{P}}$ and $\dot{\lambda}_{\mathrm{EP}}$.
(a) Trajectories.
(b) Sliding manifold and error.
(c) Steering controls $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$.
(d) $a_{\mathrm{P}}$ and $\dot{\lambda}_{\mathrm{EP}}$.
Figure 9: The defender intercepts the pursuer for various values of time $t_{2}$.
(a) Trajectories.
(b) Sliding manifold and error.
(c) Steering controls $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$.
(d) $a_{\mathrm{P}}$ and $\dot{\lambda}_{\mathrm{EP}}$.
Figure 10: The pursuer is using different guidance strategy to capture the evader.
4.2 Information-level Cooperation between the Evader and Defender
In this section, we demonstrate the effectiveness of the proposed guidance strategy when the evader and defender employ information-level cooperation to capture the pursuer. The design parameters are chosen as: $\mathcal{K}_{1}=1$, $t_{2}=4\,$\mathrm{S}$\$, $k_{3}=3$, $t_{1}=5\,$\mathrm{s}$\$, $t^{\star}=9\,$\mathrm{s}$\$, $k_{1}=k_{2}=3$, and $\mathcal{K}=1$.
Under the guidance commands derived in Theorems 5 and 4, the performance of the information-level cooperation strategy is evaluated for different values of the desired angle $\chi^{\star}$ while maintaining the same initial engagement geometry. The values of $\chi^{\star}$ are chosen to be $150^{\circ}$, $180^{\circ}$, and $200^{\circ}$, respectively and the performance is depicted in Figure 8. In this engagement scenario, the evader is initially positioned at the origin with a heading angle of $0^{\circ}$. The defender is placed at a radial distance of $500\,$\mathrm{m}$\$ from the evader with a LOS angle of $45^{\circ}$ and a heading of $45^{\circ}$, while the pursuer starts at a radial distance of $5000\,$\mathrm{m}$\$ from the evader with a heading of $180^{\circ}$, that is, directly pointing towards the evader. It can be observed from Figure 8(a) that the defender successfully captures the pursuer for all values of $\chi^{\star}$. As shown in Figures 8(c) and 8(d), the evader nullifies its LOS rate with respect to the pursuer exactly at $t_{2}=4\,$\mathrm{s}$\$ by applying the guidance command given in (31). In other words, within $4\,$\mathrm{s}$\$, the evader becomes non-maneuvering relative to the pursuer, effectively luring it by acting as bait. This behavior is further evident in Figure 8(c), where the evader’s guidance command reduces to zero precisely at $t=4\,$\mathrm{s}$\$ and remains there. During the same period of time, the pursuer applies a non-zero control input in an attempt to capture the evader. However, once the evader regulates its LOS rate relative to the pursuer, the situation becomes favorable for the pursuer. Believing interception to be assured, the pursuer also ceases maneuvering and maintains zero control input from $t=4\,$\mathrm{s}$\$ onward, as shown in Figure 8(d). Meanwhile, Figure 8(b) illustrates that the sliding mode is enforced on the inner-layer sliding manifold at $t_{1}=5\,$\mathrm{s}$\$, and the error $\beta$ converges to zero exactly at $t^{\star}=9\,$\mathrm{s}$\$ for all values of the desired angle $\chi^{\star}$. Consequently, the defender aligns itself with the desired course at the same instant, $t^{\star}=9\,$\mathrm{s}$\$, which ultimately leads to the successful capture of the pursuer for all three values of $\chi^{\star}$.
We further demonstrate the performance of the proposed (information-level cooperation) guidance strategy when the evader positions itself as bait at different time instants $t_{2}$. The values of $t_{2}$ are chosen as $4\,$\mathrm{s}$\$, $6\,$\mathrm{s}$\$, and $8\,$\mathrm{s}$\$, respectively. Keeping the other engagement parameters identical to the previous case, the results are illustrated in Figure 9. It can be observed from Figure 9(a) that the defender successfully captures the pursuer for all choices of $t_{2}$. Specifically, the evader nullifies its LOS rate with the pursuer exactly at the prescribed time instant $t_{2}$ by applying the guidance command in (31). For smaller values of $t_{2}$, the evader requires a larger initial control effort. However, once the LOS rate is nullified, the demand reduces to zero. In response, the pursuer initially applies a higher value guidance command, but as soon as the evader becomes non-maneuvering (precisely at $t_{2}$), it transitions to a straight-line path to capture the evader. By contrast, the defender utilizes the maneuvering information of the evader and always captures the pursuer, exhibiting a similar behavior across all cases.
Finally, we evaluate the performance of the proposed strategy when the pursuer employs different guidance laws to capture the evader. For simplicity, the engagement geometry is kept identical to the previous case. Here, we consider that the pursuer is using the three different strategies: PNG, Biased PNG, and PPG. In the case of BPNG, it is assumed that the pursuer has access to the evader’s maneuver information and incorporates a bias term into PNG to anticipate and counter the evader’s maneuvers. Under the proposed information level cooperation strategy, performance of the proposed strategy for such engagement scenarios is shown in Figure 10. As observed from Figure 10, the defender successfully intercepts the pursuer at the desired angle $\chi^{\star}=180^{\circ}$ in all cases, despite the different guidance strategies adopted by the pursuer. One may notice from Figure 10(d) that during the initial period, the pursuer applies different values of the guidance command to capture the evader. However, once the pursuer lures it onto the collision course (by nullifying its LOS rate with the pursuer to zero) at $t_{2}=4\,$\mathrm{s}$\$, the pursuer also stops maneuvering and aims to capture the evader within the shortest possible time by moving in a straight line directly towards the evader. Nonetheless, the defender, leveraging the information provided by the evader, intercepts the pursuer in all scenarios.
It is important to note that the defender achieves interception with only information-level cooperation from the evader, irrespective of the guidance strategy employed by the pursuer. This, in turn, bolsters the claim that the proposed guidance strategies are independent of the pursuer’s strategy.
5 Conclusions
@@
In this paper, we have introduced a geometric approach that addresses the challenge of pursuit-evasion scenarios involving three agents with arbitrary initial geometries. Our proposed solution, the evader-defender cooperative guidance strategy, offers an effective means to guarantee pursuit-evasion under diverse conditions. Specifically, it guarantees that the defender reaches a specific angle precisely within a predefined time, irrespective of the initial engagement geometry, thus intercepting the pursuer before it can capture the evader. A distinguishing feature of our approach is its adaptability to scenarios characterized by nonlinear dynamics, non-holonomic constraints, and large heading angle errors, which are often encountered in practical motion control situations, including aircraft defense. This paper presents a robust and practical solution for guaranteed pursuit-evasion problems, and also lays the foundation for identifying crucial conditions and configurations that facilitate successful evasion or capture. Analyzing geometrical solutions in three-dimensional settings could be an interesting future research topic.